Jane is a 3-year old female brought into the urgent care clinic by her mother. Child has been complaining of left ear pain for the past 2 days. She has had fevers with the last at 101 F; mother treated fever with ibuprofen. No cough, vomiting or diarrhea. You are working with a family practice Nurse Practitioner who diagnoses the child with left otitis media (ear infection). The NP selects amoxicillin suspension for treatment. The NP asks that you double check her dosing and then go through the patient teaching with parent regarding antibiotic treatment. Child is 15 kg. She takes no other medications. Answer questions based on the above information.

  1. The provider has selected 90 mg per kg dosing for this child; the suspension selected is 400 mg in 5 ml. What quantity would the child receive if the total 24-hour dose was to be divided in two doses and given by mouth every 12 hours.

  1. List two teaching considerations for parents when discussing amoxicillin treatment for bacterial infections. Do not include possible side effects or education regarding possible side effects for your answer.
    1. Mom returns to the clinic 5 days later reporting that low-grade fevers continue and that the child’s ear pain has not improved. Following assessment, the NP decides that the left ear otitis media has not improved and that the child will need to be started on oral augmentin (amoxicillin with clavulanic acid) suspension. When the NP leaves, the mom asks you why a different antibiotic needs to be started. Respond to the mother’s question using language she will understand.
    Answer #1

    A child weighs =15 kg

    Dose =90 mg per kg

    Total dose = 90* 15 =1350 mg

    formula

    Dose -= desired dose/ available dose * strength available

    =1350/400*5ml

    =16.9 ml delivers a dose which should be given in two doses 12 hrly

    So 16.9/2= 8.45 ml should be given twice daily.

    B.parentral teachings

    • Be sure your child finishes the full course of antibiotics to ensure a complete recovery from the infection and reduce the risk of antibiotic resistance.
    • medication may be ineffective if not taken as prescribed
    • Like any medical treatment, antibiotics have a risk of side effects. These can vary from mild to life-threatening.

    C.

    • Amoxicillin remains first line therapy for children who have not received amoxicillin within the past 30 days.
    • Amoxicillin/clavulanate is recommended if amoxicillin has been taken within the past 30 days, if concurrent purulent conjunctivitis is present, or if the child has a history of recurrent AOM unresponsive to amoxicillin.
    • posssiblity is there that the infecting organism is not sensitive to amoxicillin